# Introspective Planning: Guiding Language-Enabled Agents to Refine Their Own Uncertainty

### v2

@@ 4.3 Datasets
Mobile Manipulation: The original calibration or training dataset comprises 400 examples, while the test set includes 200 examples. We also evaluated the robustness to two kinds of distribution shifts: covariate shift and concept shift, using three additional datasets: one with 200 unambiguous instructions, another with 200 ambiguous instructions, and a third with 100 novel scenes and instructions. The original dataset follows the same distribution of different types of examples as in KnowNo (Ren et al., 2023), encompassing a range of types such as single-label, multi-label, spatially-ambiguous, unsafe, and Winograd tasks.
Safe Mobile Manipulation: We revised the original dataset to place increased emphasis on safety considerations, with 400 examples for calibration and 200 for testing. We have also assembled an additional dataset of 200 safety-centric examples. These safety-related scenarios can be categorized into three types: (1) ambiguous instructions that become clear when safety is considered, e.g., choosing between a metal and plastic bowl for microwave use, (2) ambiguous instructions considering safety, such as selecting the correct bowl for microwave heating among stainless steel, plastic, or ceramic options, and (3) unambiguous but unsafe instructions, such as ’place a metal bowl in the microwave’.
Tabletop Rearrangement: The task involves moving colored blocks and bowls on a table according to specific instructions. These instructions are intentionally designed to include ambiguities in attributes (such as alternative names for objects and colors), numbers (using vague terms for quantities), and spatial relationships (using general terms for directions and orientations). For this dataset, 400 examples were used for calibration, and an additional 200 examples were employed for testing. The experiment results on this dataset are in Appendix C and the details for each dataset are included in Appendix A.
@@ 5 Results
We implemented all tasks using OpenAI’s GPT-3.5 (text-davinci-003) and GPT-4 Turbo (gpt-4-1106-preview). Results on GPT-3.5 are presented in the main paper, with GPT-4 Turbo findings detailed in Appendix C. We employed Sentence-BERT (Reimers & Gurevych, 2019) to encode instructions, retrieving the top $m=3$ text embeddings based on cosine similarity. The size of our knowledge base is the same as the calibration set which is 400. However, experimental explorations with sizes of 10, 50, 100, and 200 indicated that a smaller knowledge base, specifically of 100 or 200, effectively achieves satisfactory results. The relationship between performance and knowledge base size is further detailed in Appendix B.
Table 1: GPT-3.5 Results for Mobile Manipulation. Original: Original mobile manipulation dataset. All Unambiguous: Datasets contain only unambiguous instructions. All Ambiguous: Datasets contain only ambiguous instructions. Novel Data: Noval data that contains unseen objects and instructions. SR: Success rate, HR: Help rate, ESR: Exact Set Rate. Note that only KnowNo uses conformal prediction. The target success rate for KnowNo is $85\%$. All numbers are in percentages.
Original All Unambiguous All Ambiguous Novel Data
Method SR $\uparrow$ HR $\downarrow$ ESR $\uparrow$ SR $\uparrow$ HR $\downarrow$ ESR $\uparrow$ SR $\uparrow$ HR $\downarrow$ ESR $\uparrow$ SR $\uparrow$ HR $\downarrow$ ESR $\uparrow$
KnowNo (w. CP) 83.0 74.5 25.0 90.5 67.0 31.5 84.5 86.5 28.0 80.0 86.0 28.0
Prompt Set 77.0 70.0 44.5 81.0 36.0 51.0 80.0 90.5 52.5 73.0 64.0 46.0
Prompt Set + CoT 83.5 72.5 50.5 79.5 32.5 48.5 80.5 92.0 55.5 75.0 69.0 52.0
Retrieval-Q-CoT 73.0 56.0 50.0 78.0 12.0 70.0 71.0 74.0 41.0 68.0 56.0 44.0
Auto-CoT 71.5 41.0 47.5 78.5 2.0 70.0 72.0 60.5 35.0 67.0 43.0 49.0
Ours (w.o CP) 92.0 69.0 78.5 91.0 14.5 86.0 94.0 97.5 79.5 88.0 70.0 73.0
@@
Table 2: GPT-3.5 Results for Safe Mobile Manipulation. Original + Safety Data: Original mobile manipulation dataset with additional safety-related data. Only Safety Data: Dataset that contains only safety-related data. SR: Success rate, HR: Help rate, ESR: Exact Set Rate, NCR: Noncompliance contamination rate, UCR: Unsafe contamination rate. Note that only KnowNo uses conformal prediction. The target success rate for KnowNo is $85\%$. All numbers are in percentages.
Original + Safety Data Only Safety Data
Method SR $\uparrow$ HR $\downarrow$ ESR $\uparrow$ NCR $\downarrow$ UCR $\downarrow$ SR $\uparrow$ HR $\downarrow$ ESR $\uparrow$ NCR $\downarrow$ UCR $\downarrow$
KnowNo (w. CP) 84.5 94.0 10.0 83.5 12.0 87.0 98.0 7.0 50.0 85.0
Prompt Set 69.5 57 47.0 28.5 11.5 35.0 34.0 15.0 6.0 85.0
Prompt Set + CoT 69.5 63.0 55.5 24.5 6.5 55.0 53.0 50.0 15.0 34.0
Retrieval-Q-CoT 66.0 54.0 47.0 24.0 7.5 51.0 47.0 39.0 17.0 50.0
Auto-CoT 44.0 38.0 25.0 37.5 7.5 36.0 32.0 35.0 20.0 59.0
Ours (w.o CP) 91.0 69.0 80.5 12.0 1.0 96.0 76.0 93.0 2.0 5.0
(a) Results on Mobile manipulation dataset.
(b) Results on Safe mobile manipulation dataset
Figure 4: Success Rate vs. Target Success Rate using conformal prediction on both dataset. We found that conformal prediction can provide a statistical guarantee of task completion for all three methods.
(a) ESR vs. TSR
(b) HR vs. TSR
(c) Average Set Size vs. TSR
(d) Confidence Quantile vs. TSR
Figure 5: We evaluate exact set rate, help rate, average prediction set size, and confidence quantile of each models using conformal prediction with Mobile Manipulation Dataset from KnowNo (Ren et al., 2023)
(a) ESR vs. TSR
(b) HR vs. TSR
(c) Average Set Size vs. TSR
(d) Confidence Quantile vs. TSR
Figure 6: We evaluate exact set rate, help rate, average prediction set size, and confidence quantile of each models using conformal prediction with Safe Mobile Manipulation Dataset.
(a) Noncompliant Contamination Rate vs. TSR
(b) Unsafe Contamination Rate vs. TSR
Figure 7: We evaluate Noncompliant Contamination Rate and Unsafe Contamination Rate using conformal prediction on Safe Mobile Manipulation Dataset.
@@ 5.1 Direct Prediction
In Table 1, our analysis highlights that introspective planning with direct prediction outperforms all baseline methods in Success Rate (SR) and Exact Set Rate (ESR), indicating superior predictive accuracy. Retrieve-Q-CoT and Auto-CoT don’t outperform other prompt methods on the original dataset, since the model may potentially generate misleading reasoning in the training phase without access to ground truth actions. However, they excel in the unambiguous instruction dataset, achieving 70% ESR, despite lower success and minimal help rates across all datasets. This observed trend suggests both methods lean towards generating incorrect predictions with high confidence, thereby highlighting the critical role of introspective planning in offering meaningful guidance for enhancing LLM reasoning capabilities. Moreover, as shown in Table 2, our approach significantly outperforms all baselines in safety-critical scenarios, exhibiting impressive improvements in safety prioritization and dramatically reducing Noncompliance Contamination Rate (NCR) and Unsafe Contamination Rate (UCR).
Through additional conformal prediction, KnowNo is able to achieve a notably higher success rate in comparison to Prompt Set at the sacrifice of low Exact Set Rate (ESR)—indicating the inclusion of incorrect options. In contrast, the direct prediction method is distinguished by its ability to generate more precise and accurate prediction sets, with the Prompt Set method substantially improving the ESR by nearly 19.5% over KnowNo, an enhancement that is further augmented through the strategic integration of Chain of Thought (CoT) techniques. In the next section, we’ll delve into the advantages of integrating introspective planning with conformal prediction, demonstrating their superiority over traditional prompting methods in detail.
@@ 5.2 Conformal Prediction
In section 3, we compare introspective planning with two baselines using conformal prediction, assessing the influence of Target Success Rate (target success rate) through ablation studies. The findings (3(a), 7(b)) confirm that conformal prediction aligns the empirical success rate with target success rate in both datasets. Notably, 4(a) shows that our method consistently achieves a higher Exact Set Rate across the full range of target success rate, outperforming the Retrieval-Q-CoT, which only shows marginal improvement over KnowNo at target success rate levels below $85\%$. This superiority is echoed in the safe manipulation dataset (5(a)), validating our approach’s robustness.
Our analysis indicates Retrieve-Q-CoT and KnowNo are less certain about correct choices, leading to larger prediction sets for achieving desired success rate levels. To validate this observation, we calculated the average prediction set size of each method and have detailed our findings in 4(c) and 5(c). Our analysis revealed that introspective planning tends to include fewer options in its prediction sets for a given target success rate. This is also reflected in each model’s help rate across both datasets, as illustrated in 4(b) and 5(b), where it is evident that both KnowNo and Retrieval-Q-CoT require human clarifications more frequently than introspective planning. From the conformal prediction perspective, we noticed that introspective planning has a higher calibration threshold $1-\hat{q}$ than two baseline methods in both datasets (4(d) and 5(d)), which indicates our methods achieve a tighter confidence bound for the statistical guarantee. These comprehensive empirical results strongly suggest that introspective planning offers a more precise reasoning framework when navigating uncertainties in robot planning scenarios.
Furthermore, we also study whether introspective planning helps to better reason about task ambiguities and unsafe actions in robot planning. Preliminary results from the Safe Mobile Manipulation dataset show introspective planning significantly reduces non-compliant contamination rates compared to baseline methods and maintains a lower unsafe contamination rate across all target success rate levels. Additionally, we demonstrate how the conformal prediction framework can be extended for multi-label predictions, with performance details in Appendix D.
6 Related Works
LLM as a reasoner. Known as zero-shot chain-of-thought (Kojima et al., 2022), LLMs can generate multiple reasoning steps by supplying a simple prompt ”Let’s think step by step”. The Chain-of-thought (CoT) reasoning performance is further improved by providing LLMs with manually designed examples (Wei et al., 2022b). Tree-of-Thoughts (Yao et al., 2023) further generalizes CoT by considering multiple reasoning paths. The most relevant prior work is Auto-CoT (Zhang et al., 2023), which leverages zero-shot CoT to generate diverse reasoning chains as demonstrations and samples them as examples for in-context learning. However, it does not use ground truth to generate knowledge, leading to mispredictions from noisy thoughts.
LLM as a planner. Emergent reasoning allows LLMs to break down a task into intermediate subgoals and generate actions as a sequence of plans (Huang et al., 2022a). Through prompting and in-context learning, LLMs can ground human instructions in natural language into executable robot actions conditioned on scene descriptions (Liang et al., 2023; Singh et al., 2023; Ahn et al., 2022; Huang et al., 2022b). Recent works further enhance the reasoning and planning ability by iteratively refining actions through self-reflection during planning (Madaan et al., 2023; Paul et al., 2023; Chang, 2023). The planning performance can also benefit from summarizing previous failures as feedback to guide future planning (Shinn et al., 2023; Wang et al., 2023; Liu et al., 2023). However, rather than relying on feedback from past actions, it uses demonstrations to guide the LLM’s reasoning about safety and uncertainty without the need for iterative refinement.
Quantifying uncertainty in LLMs. Understanding uncertainty in LLMs is essential to validate the robustness of task planning and identify potential failure cases. Growing interest from the natural language processing community has spurred efforts to quantify uncertainty in LLM generation (Ott et al., 2018; Fomicheva et al., 2020; Xiao et al., 2022; Baan et al., 2023), calibrate uncertainty (Desai & Durrett, 2020; Jiang et al., 2021; Zhao et al., 2021; Yin et al., 2023), and examine LLM reliability (Ji et al., 2023b; Lin et al., 2021). KnowNo (Ren et al., 2023) first tackles the uncertainty quantification and calibration in LLM planners by casting planning as multiple choice question answering (MCQA). It uses conformal prediction to output a set of candidate plans with a calibrated target probability of containing a successful one. Our conformal prediction evaluation follows the same framework as KnowNo, and our extended analysis of planning performance identifies potential directions for future improvement.
@@ 7 Conclusion
In this paper, we investigate introspective planning, a systematic guidance process that allows LLMs to account for uncertainty and safety in their plan generation. We conducted thorough evaluations on three different datasets and found that introspective planning improves upon the state of the art across all metrics. In addition, we show that introspection can be integrated with the conformal prediction framework, achieving a tighter confidence bound and maintaining statistical success guarantees with fewer superfluous user clarification queries.
Limitation. While the proposed introspective reasoning method yields significant prediction improvements, we acknowledge two major limitations in quantifying uncertainties and guiding LLM agents. First, as our method is designed and calibrated to mirror the exact set of valid actions, it fails to differentiate between over-confident and under-confident predictions. Second, though we demonstrated lower unsafe and non-compliant contamination with our approach, the quantification method does not adequately distinguish practical differentiation between safety and compliant uncertainty. One potential direction for future works is to quantify coverage and purity simultaneously through multi-label prediction with conformal calibration.
@@ 8 Broader Impact
Appendix A Details of the dataset

### v3

@@ 4.2 Datasets
Mobile Manipulation: The original calibration or training dataset comprises 400 examples, while the test set includes 200 examples. We also evaluated the robustness to two kinds of distribution shifts: covariate shift and concept shift, using three additional datasets: one with 200 unambiguous instructions, another with 200 ambiguous instructions, and a third with 100 novel scenes and instructions. The original dataset follows the same distribution of different types of examples as in KnowNo [31], encompassing a range of types such as single-label, multi-label, spatially-ambiguous, unsafe, and Winograd tasks. The experiment results on this dataset are in Appendix A.
Safe Mobile Manipulation: We augment the original mobile manipulation dataset to emphasize safety, with 400 examples for calibration and 200 for testing. Additionally, we assembled a dataset of 200 safety-critical scenarios, categorized into three types: (1) ambiguous instructions that become clear when safety is considered, e.g., choosing between a metal and plastic bowl for microwave use, (2) ambiguous instructions considering safety, such as selecting the correct bowl for microwave heating among stainless steel, plastic, or ceramic options, and (3) unambiguous but unsafe instructions, such as ‘place a metal bowl in the microwave’.
Tabletop Rearrangement: The task involves moving colored blocks and bowls on a table according to specific instructions. These instructions are intentionally designed to include ambiguities in attributes (such as alternative names for objects and colors), numbers (using vague terms for quantities), and spatial relationships (using general terms for directions and orientations). For this dataset, 400 examples were used for calibration and an additional 200 examples for testing. The experiment results on this dataset are in Appendix A andeach dataset are detailed in Appendix B.
@@ 5 Results
Table 1: GPT-4 Results for Safe Mobile Manipulation. SR: Success rate, HR: Help rate, OAR: Over-Ask rate, OSR: Over-Step rate, UR: Unsafe rate, ESR: Exact Set Rate, NCR: Noncompliance contamination rate, UCR: Unsafe contamination rate. Conformal means conformal prediction and Direct means direct prediction. All the others use direct prediction. The target success rate for conformal prediction is $85\%$. All numbers are reported in percentages.
Decision Metrics System Metrics
Method ESR $\uparrow$ NCR $\downarrow$ UCR $\downarrow$ SR $\uparrow$ HR OAR $\downarrow$ OSR $\downarrow$ UR $\downarrow$
KnowNo (Conformal) 37.5 51.0 7.0 84.5 77.5 51.3 35.3 1.0
Prompt Set 73.5 11.5 3.5 82.5 63.0 3.8 36.2 2.5
Prompt Set + CoT 79.0 10.0 5.0 87.5 67.0 10.3 30.8 4.0
Retrieval-Q-CoT 81.5 7.0 4.5 88.0 65.0 2.6 26.1 4.0
Auto-CoT 77.5 10.0 5.0 85.5 62.5 1.3 37.3 4.0
Ours (Conformal) 58.0 27.5 3.0 87.5 63.0 6.4 21.6 1.5
Ours (Direct) 93.0 5.5 0.5 96.5 67.5 0.0 3.8 0.5
Implementation details. We implemented all tasks using OpenAI’s GPT-3.5 (text-davinci-003) and GPT-4 Turbo (gpt-4-1106-preview). We only presented GPT-4 Turbo results on Safe Mobile Manipulation in the main paper. All of the other results are in Appendix A. We employed Sentence-BERT [30] to encode instructions, retrieving the top $m=3$ text embeddings based on cosine similarity. We used the default temperature of 0 to sample the LLM’s response. The knowledge base and the calibration set contain 400 tasks each. Appendix C contains an in-depth experimental exploration of performance variation with knowledge base size, with sizes of 10, 50, 100, and 200, suggesting that modest knowledge bases with around 100 examples still enable satisfactory results.
Trade-off between direct and conformal prediction: Our experiments indicate that introspective planning with direct prediction significantly outperforms all other baselines in terms of performance metrics. However, this approach does not guarantee success. On the other hand, introspective planning with conformal prediction guarantees success and surpasses other methods employing conformal prediction. Nevertheless, a noticeable performance gap exists between direct prediction and conformal prediction. This highlights an intriguing trade-off: while conformal prediction provides success guarantees, it tends to be more conservative.
Figure 4: Variation of different performance metrics with respect to the Target Success Rate (TSR). Each subplot compares KnowNo, Retrieval-Q-CoT, and Ours (Conformal) methods across various metrics. Introspective planning (Ours-Conformal) consistenty achieves the best tradeoff between performance metrics and Target Success Rate (TSR) across all comparisons.
@@ 5.1 Direct Prediction
In Tab. 1, our analysis highlights that introspective planning with direct prediction outperforms all baseline methods in both decision and prediction metrics. KnowNo is too conservative with low and Non-compliant contamination rate (NCR) and exact set rate (ESR). As a result, despite guaranteeing a high success rate, it suffers from a very high Over-Ask Rate (OAR), which is not ideal. Without conformal prediction, the Prompt Set method generates more accurate prediction sets but sacrifices the guarantee of success, as indicated by the 19.5% increase in ESR. Using Chain of Thought (CoT) further improves performance. Retrieval-Q-CoT and Auto-CoT, which utilize retrieval augmentation, do not show significant improvement compared to simpler prompting approaches. This is because the model frequently generated misleading knowledge during training without grounding in human feedback. Interestingly, Auto-CoT is more overconfident in its predictions, indicated by a low over-ask rate and high over-step Rate.
Compared to other baselines, Introspective planning guides the LLM to generate more precise prediction sets, as evidenced by the highest exact set rate and lowest non-compliant contamination rate. It avoids over-asking, rarely oversteps, and has the lowest unsafe contamination rate and unsafe rate, demonstrating effective reasoning about both uncertainty and safety.
@@ 5.2 Conformal Prediction
In Figure 4, we compare introspective planning with two baselines using conformal prediction. The findings (Figure 4(a)) confirm that conformal prediction aligns the empirical success rate with target success rate. Notably, Figure 4(d) shows that our method consistently achieves a higher Exact Set Rate across the full range of target success rate, outperforming both KnowNo and the Retrieval-Q-CoT.
Our analysis indicates that both Retrieve-Q-CoT and KnowNo are more conservative and generate larger prediction sets to achieve desired success rate levels, as indicated by Figure 4(c). Consequently, they more frequently include irrelevant options, resulting in high non-compliant contamination rate (NCR) as shown in Figure 4(e). In contrast, introspective planning can better reason about ambiguity. It excels in unambiguous scenarios, as both Retrieval-Q-CoT and KnowNo over-ask much more frequently than introspective planning across the target success rate, as indicated by Figure 4(g). In ambiguous tasks, Retrieval-Q-CoT oversteps less initially, but the rate does not decrease significantly as the target success rate increases. Conversely, our approach effectively reduces the overstepping rate while maintaining the lowest over-asking rate, as shown in Figure 4(h). Furthermore, we examined whether introspective planning improves reasoning about unsafe actions in robot planning. Results in Figure 4(f) show that introspective planning maintains the lowest Unsafe Contamination Rate (UCR) across all target success rate levels, indicating its effectiveness in reasoning about unsafe options.
From the conformal prediction perspective, we observed that introspective planning has a lower $\hat{q}$, resulting in a higher calibration threshold $1-\hat{q}$, compared to the two baselines, as shown in Figure 4(i). This indicates that our method achieves a tighter confidence bound for the statistical guarantee.
6 Related Work
LLMs as reasoning engines. Through a process known as (zero-shot) chain-of-thought (CoT), LLMs can be prompted to generate multiple reasoning steps by instructing them to “think step by step” at inference time [18]. This method’s accuracy can be improved by including manually designed examples (few-shot CoT) [41]. The tree-of-thoughts approach [44] generalizes CoT by considering multiple reasoning paths. Our work is inspired by the retrieval augmentation mechanism in Auto-CoT [49] and Retrieval-Q-CoT [49], which first use zero-shot CoT to generate a diverse set of reasoning chains and then sample them at runtime as few-shot CoT examples. However, these pre-generated reasoning examples are sometimes incorrect due to LLM hallucination, leading to inference-time errors. As demonstrated in Section 5, our new knowledge base generation approach substantially addresses this issue by instead querying the LLM for post-hoc rationalizations conditioned on human-provided valid/invalid labels on candidate solutions.
Retrieval-augmented generation. Retrieval-augmented generation (RAG) augments the input space of LMs with retrieved text passages, significantly improving performance on knowledge-intensive tasks [10, 20, 29]. Traditional RAG methods typically use open-source, off-the-shelf knowledge bases for text generation [16, 33, 25, 47, 43, 51]. Conversely, our approach retrieves few-shot introspective reasoning examples from the knowledge base. This guides LLMs to explicitly reason about uncertainties and safety, formulating plan in a structured format, as shown in Tab. 13. In practice, we observe this strategy results in a more grounded reasoning process compared to conventional RAG, which relies on open-source knowledge bases. While existing RAG literature primarily addresses content hallucination, our approach aims to equip language agents with the capability to introspect and refine their own uncertainties. This emphasis allows for uncertainty-aware planning in robotics and achieves tighter statistical guarantees with conformal prediction.
LLMs as planners. Emergent reasoning allows LLMs to break down a task into intermediate subgoals and generate actions as a sequence of plans [11]. Through prompting and in-context learning, LLMs can ground human instructions in natural language into executable robot actions conditioned on scene descriptions [21, 35, 1, 12]. Recent works further enhance the reasoning and planning ability by iteratively refining actions through self-reflection during planning [45, 24, 28, 6, 34, 39, 23]. ReAct [45] and Reflexion [34] focus on multi-step planning scenarios, in which robots can execute certain actions, observe the state feedback, and replan for correction. However, in safety-critical robotic applications, certain invalid actions can immediately lead to catastrophic safety failures that cannot be recovered from. Therefore, instead of relying on self-correction by trial and error, our method uses retrieval augmentation to guide the language agent to proactively reason about task compliance and safety at the planning stage. Additionally, recent work [17] has shown that LLMs cannot plan effectively through self-reflection alone but can do so when integrated with external verifiers, aligning with our method that employs LLMs to support planning by constructing an external knowledge base.
Quantifying uncertainty in LLMs. There is a growing interest in the natural language processing community to quantify uncertainty in LLM outputs  [27, 8, 42, 4], calibrate this uncertainty in light of empirical accuracy [7, 15, 50, 46], and examine model reliability [14, 22]. Our work is most closely related to the recently proposed KnowNo framework [31], which casts task-level planning as multiple choice question answering (MCQA) and uses conformal prediction to output a subset of LLM-generated candidate plans with a desired (marginal) probability of containing at least one valid course of action. Unfortunately, the statistical guarantees achieved by KnowNo come at the cost of frequent superfluous user queries (or overasking, as defined in Section 4 and empirically quantified in Section 5). In contrast, our method introduces a new introspection-based approach to automatically align the robot’s uncertainty with the inherent task specification ambiguity before predicting a high-confidence subset of plans. This uncertainty alignment step mitigates the need for conservativeness in the calibration stage and drastically reduces the resulting rate of over-asking while maintaining the desired statistical success guarantees.
@@ 7 Conclusion
This paper proposes and investigates a novel introspective planning framework that allows language-enabled agents to align their decision-making uncertainty with safety and task ambiguity. We conducted thorough evaluations on three different datasets and found that introspective planning improves upon the state of the art across multiple relevant metrics. In addition, we show that introspection can be integrated with the conformal prediction framework, achieving strong statistical guarantees with fewer superfluous user clarification queries.
Limitation. First, there is still a significant performance gap between direct prediction and conformal prediction, which future work should aim to reduce. Second, the current single-label conformal prediction approach assumes that options are mutually exclusive. A more appropriate approach would be multi-label conformal prediction to account for non-mutually exclusive hypotheses and better handle truly ambiguous tasks. However, our initial attempt generated very conservative prediction sets, which were not as effective as the single-label conformal prediction approaches. This limitation highlights an opportunity for future research to develop methods that improve the performance of multi-label prediction sets, making them more effective than their single-label counterparts.
@@ 8 Broader Impact
Appendix A Additional Quantitative Results
Table 2: GPT-3.5 Results for Mobile Manipulation. SR: Success rate, HR: Help rate, OAR: Overask rate, OSR: Overstep rate, ESR: Exact Set Rate, NCR: Noncompliant contamination rate. Conformal means conformal prediction and Direct means direct prediction. The target success rate for conformal prediction is $85\%$. All the others use direct prediction. All numbers are in percentages.
Decision Metrics Prediction Metrics
Method SR $\uparrow$ HR OAR $\downarrow$ OSR $\downarrow$ ESR $\uparrow$ NCR $\downarrow$
KnowNo (Conformal) 83.0 74.5 61.9 58.8 25.0 47.0
Prompt Set 77.0 70.0 34.9 46.7 44.5 35.0
Prompt Set + CoT 83.5 72.5 28.6 34.5 50.5 32.5
Retrieval-Q-CoT 73.0 56.0 11.1 54.3 50.0 21.5
Auto-CoT 71.5 41.0 4.76 66.1 47.5 20.0
Ours (Conformal) 85.0 59.5 11.1 36.6 51.5 26.5
Ours (Direct) 92.0 69.0 3.17 11.1 78.5 18.5
Table 3: GPT-3.5 Results for Safe Mobile Manipulation. SR: Success rate, HR: Help rate, OAR: Overask rate, OSR: Overstep rate, UR: Unsafe rate, ESR: Exact Set Rate, NCR: Noncompliance contamination rate, UCR: Unsafe contamination rate. Conformal means conformal prediction and Direct means direct prediction. The target success rate for conformal prediction is $85\%$. All the others use direct prediction. All numbers are reported in percentages.
Decision Metrics Prediction Metrics
Method SR $\uparrow$ HR OAR $\downarrow$ OSR $\downarrow$ UR $\downarrow$ ESR $\uparrow$ NCR $\downarrow$ UCR $\downarrow$
KnowNo (Conformal) 84.5 94.0 88.5 27.3 0.5 10.0 83.5 12.0
Prompt Set 69.5 57.0 20.5 57.3 5.5 47.0 28.5 11.5
Prompt Set + CoT 69.5 63.0 6.4 48.3 4.5 55.5 24.5 6.5
Retrieval-Q-CoT 66.0 54.0 3.8 55.4 6.5 47.0 24.0 7.5
Auto-CoT 44.0 38.0 1.3 75.4 7.0 25.0 37.5 7.5
Ours (Conformal) 85.0 71.5 9.0 26.3 0.0 51.0 37.0 0.0
Ours (Direct) 91.0 69.0 0.0 16.7 0.5 80.0 12.0 1.0
Table 4: GPT-3.5 Results for further studies on Mobile Manipulation. All Unambiguous: Datasets contain only unambiguous instructions. All Ambiguous: Datasets contain only ambiguous instructions. Novel Data: Novel data that contains unseen objects and instructions. SR: Success rate, HR: Help rate, ESR: Exact Set Rate. Conformal means conformal prediction and Direct means direct prediction. The target success rate for conformal prediction is $85\%$. All the others use direct prediction. All numbers are reported in percentages.
All Unambiguous All Ambiguous Novel Data
Method SR $\uparrow$ HR ESR $\uparrow$ SR $\uparrow$ HR ESR $\uparrow$ SR $\uparrow$ HR ESR $\uparrow$
KnowNo (Conformal) 90.5 67.0 31.5 84.5 86.5 28.0 80.0 86.0 28.0
Prompt Set 81.0 36.0 51.0 80.0 90.5 52.5 73.0 64.0 46.0
Prompt Set + CoT 79.5 32.5 48.5 80.5 92.0 55.5 75.0 69.0 52.0
Retrieval-Q-CoT 78.0 12.0 70.0 71.0 74.0 41.0 68.0 56.0 44.0
Auto-CoT 78.5 2.0 70.0 72.0 60.5 35.0 67.0 43.0 49.0
Ours (Direct) 91.0 14.5 86.0 94.0 97.5 79.5 88.0 70.0 73.0
Analysis on GPT-3.5 We evaluated our method on Mobile Manipulation and Safe Mobile Manipulation using GPT-3.5. The outcomes are presented in Tab. 2 and Tab. 3. Introspective planning still demonstrates the strongest capability to reason about uncertainty and safety, as indicated by the highest success rate and exact set rate, as well as the lowest non-compliance rate, unsafe contamination rate. It also overasks and oversteps much less than all the other approaches. While GPT-3.5 is relatively weaker compared to GPT-4, the positive impact of introspective planning is even more pronounced. We also conducted an analysis of two types of distribution shifts: covariate shift and concept shift. IntroPlan performs the best in both datasets: one with only unambiguous scenarios and one with only ambiguous scenarios. Additionally, it effectively generalizes to novel scenes, as indicated by a significantly higher exact set rate.
Table 5: GPT-4 Results for Mobile Manipulation. SR: Success rate, HR: Help rate, OAR: Overask rate, OSR: Overstep rate, ESR: Exact Set Rate, NCR: Noncompliant contamination rate. Conformal means conformal prediction and Direct means direct prediction. The target success rate for conformal prediction is $85\%$. All the others use direct prediction. All numbers are reported in percentages.
Prediction Metrics Decision Metrics
Method ESR $\uparrow$ NCR $\downarrow$ SR $\uparrow$ HR OA $\downarrow$ OS $\downarrow$
KnowNo (Conformal) 37.0 45.5 83.0 72.0 47.6 37.5
Prompt Set 76.0 16.0 90.0 65.5 11.1 26.1
Prompt Set + CoT 81.0 11.0 93.0 72.5 23.8 7.4
Retrieval-Q-CoT 84.0 14.0 96.5 70.5 11.1 12.3
Auto-CoT 81.0 16.5 97.0 72.0 23.8 19.3
Ours (Conformal) 53.5 34.5 83.0 60.5 3.2 25.3
Ours (Direct) 94.5 9.0 99.5 68.5 3.2 3.1
Table 6: GPT-4 Results for further studies on Mobile Manipulation. All Unambiguous: Datasets contain only unambiguous instructions. All Ambiguous: Datasets contain only ambiguous instructions. Novel Data: Noval data that contains unseen objects and instructions. SR: Success rate, HR: Help rate, ESR: Exact Success Set Prediction Rate. Conformal means conformal prediction and Direct means direct prediction. The target success rate for conformal prediction is $85\%$. All the others use direct prediction. All numbers are reported in percentages.
All Unambiguous All Ambiguous Novel Data
Method SR $\uparrow$ HR ESR $\uparrow$ SR $\uparrow$ HR ESR $\uparrow$ SR $\uparrow$ HR ESR $\uparrow$
KnowNo (Conformal) 95.0 36.5 61.0 93.5 85.5 45.5 85.0 61.0 37.0
Prompt Set 81.0 36.0 51.0 80.0 90.5 52.5 73.0 64.0 46.0
Prompt Set + CoT 87.5 16.5 74.5 91.0 89.0 74.0 87.0 72.0 68.0
Retrieval-Q-CoT 86.5 8.5 82.0 93.5 88.5 75.5 91.0 66.0 76.0
Auto-CoT 93.5 23.5 71.0 93.5 89.0 79.5 91.5 73.0 75.0
Ours (Direct) 98.5 1.0 96.5 98.5 90.5 90.0 92.0 65.0 84.0
Table 7: GPT-4 Results on Safe Mobile Manipulation with only safety-critical scenarios. SR: Success rate, HR: Help rate, ESR: Exact Success Set Prediction Rate, NCR: Noncompliance contamination rate, UCR: Unsafe contamination rate. Conformal means conformal prediction and Direct means direct prediction. The target success rate for conformal prediction is $85\%$. All the others use direct prediction. All numbers are reported in percentages.
Only Safety-Critical Data
Method SR $\uparrow$ HR ESR $\uparrow$ NCR $\downarrow$ UCR $\downarrow$
KnowNo (Conformal) 83.0 78.0 22.0 28.0 64.0
Prompt Set 55.0 43.0 54.0 12.0 31.0
Prompt Set + CoT 47.0 30.0 44.0 17.0 54.0
Retrieval-Q-CoT 55.0 37.0 54.0 17.0 40.0
Auto-CoT 45.0 29.0 42.0 15.0 52.0
Ours (Direct) 93.0 72.0 90.0 6.0 6.0
Analysis on GPT-4: In the main paper, we discussed our evaluation results on Safe Mobile Manipulation with GPT-4. Here, we provide additional evaluation on Mobile Manipulation. As shown in Tab. 5, our approach consistently outperforms all other baselines. It is also the most effective in handling distribution shifts, as indicated in Tab. 6. Additionally, we evaluated Safe Mobile Manipulation in safety-critical scenarios to assess the planner’s ability to prioritize safety in Tab. 7. IntroPlan demonstrated effective task compliance, as evidenced by a high exact set rate and low non-compliant contamination rate, and ensured safety with a low unsafe contamination rate.
Figure 5: Variation of different performance metrics with respect to the Target Success Rate on Mobile Manipulation using GPT-3.5. Each subplot compares KnowNo, Retrieval-Q-CoT, and Ours (Conformal) methods across various metrics. Introspective planning (Ours-Conformal) consistently achieves the best tradeoff between metrics and Target Success Rate across all comparisons.
Figure 6: Variation of different performance metrics with respect to the Target Success Rate on Safe Mobile Manipulation using GPT-3.5. Each subplot compares KnowNo, Retrieval-Q-CoT, and Ours (Conformal) methods across various metrics. Introspective planning (Ours-Conformal) consistently achieves the best tradeoff between metrics and Target Success Rate across all comparisons.
(a)              Mobile Manipulation
(b)          Safe Mobile Manipulation
(c)          Tabletop Rearrangement
Figure 7: Variation of different performance metrics with respect to the Target Success Rate on three datasets Mobile Manipulation, Safe Mobile Manipulation, and Tabletop Rearrangement using GPT-4. Each subplot compares KnowNo, Retrieval-Q-CoT, and Ours (Conformal) methods across various metrics. Introspective planning (Ours-Conformal) consistently achieves the best tradeoff between performance metrics and Target Success Rate across all comparisons. It guarantees success, provides the most accurate prediction set, and achieves the tightest guarantee bound.
Appendix B Details of the dataset